# Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights

### v2

@@ 2. This follows from the inequality ($n\in\mathbb{N}_{0}$)
Propositions 2.19 for $n=2$ and item 1 of 2.25 are incomparable. The former assumes only conditional convergence, but for all three series. The latter assumes stronger absolute convergence, but only for two series. Another variant of the proposition is the theorem due to Franz (Franciszek) Mertens (1840–1927).
@@
Let $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=V(x)W(x)$ and if $V(1)_{1}$ and $W(1)_{2}$ exist, then $U(1)_{1}$ exists and
@@
$S_{m}=\prod_{i=1}^{m}(v_{1}+v_{2}+\dots)=\sum_{\langle j_{1},\,\dots,\,j_{m}\rangle\in\mathbb{N}^{m}}v_{j_{1}}\cdot\ldots\cdot v_{j_{m}}$}
By Proposition 2.6 it absolutely converges and has the sum $(V(1)_{2}-1)^{m}$. Thus by Propositions 2.2 and 2.4, $S$ has the sum $1+\sum_{m\geq 1}(-1)^{m}(V(1)_{2}-1)^{m}=\frac{1}{V(1)_{2}}$.
@@
By Proposition 2.4, $S$ has also the sum $u_{0}+u_{1}+u_{2}+\cdots=U(1)_{2}$ and the stated formula follows.
@@ 2. Indeed,
By Proposition 2.4, $S$ has the sum $v_{0}+v_{1}+\dots=V(1)_{2}$. By item 1 of Proposition 2.25, $U(1)_{2}$ exists and $U(1)_{2}=\big(V(1)_{2}\big)^{2}$. $\Box$
@@
Finally, we consider light convex weights.
@@
We can add and remove the condition $\ell(w^{\prime})\in V(h)$ without affecting the sum because walks $w^{\prime}$ not satisfying it have zero weights. $\Box$
@@ 3 Extensions of Pólya’s theorem with $v=1$
In this section we establish the first group of generalizations of Pólya’s theorem, namely Theorems 3.23.7. We treat the case $v=1$ when the vertex to be visited by walks coincides with their starting vertex.
We introduce some generating functions in $\mathbb{C}[[x]]$. Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ is the set of walks that visit $v$ at a step $i>0$. In this section, $v=1$. We are interested in the generating function
$A_{h}(x)=\sum_{n\geq 0}a_{n}^{h}x^{n}=\sum_{n\geq 0}h(W(1,\,n))x^{n}\,.$}
We also define
@@
where $W_{b}(n)$ ($\subset W(n)$) are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and $W_{c}(n)$ ($\subset W_{b}(n)$) are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$ but avoid $1$ between. The coefficients of $A_{h}(x)$, $\dots$, $D_{h}(x)$ are correctly defined because they are sums of subseries of absolutely convergents series. We have $a_{0}^{h}=c_{0}^{h}=0$ and $b_{0}^{h}=d_{0}^{h}=1$.
@@ 2. We show that
$F(w)=\langle w_{1},\,w_{2},\,\dots,\,w_{j}\rangle\ \ (\in W_{c}(n_{i}),\,n_{i}=m_{i}-m_{i-1})$
@@
For better orientation we label our theorems by the triples
@@
where “gen” refers to general edge weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and “con” to convex ones. The generating functions $A_{h}(x)$, $\dots$, $D_{h}(x)$ are as above
@@
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Suppose that $D_{h}(1)_{2}$ and $C_{h}(1)_{2}$ exist, and that $\sum_{n\geq 1}|b_{n}^{h}|<1$. Then $B_{h}(1)_{2}$ exists and is nonzero, $A_{h}(1)_{2}$ exists and
@@
Proof. By item 1 of Proposition 3.1 we have $A_{h}(x)=C_{h}(x)D_{h}(x)$. By item 1 of Proposition 2.25, $A_{h}(1)_{2}$ exists and $A_{h}(1)_{2}=C_{h}(1)_{2}D_{h}(1)_{2}$ . By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. By item 1 of Proposition 2.27, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h}(1)_{2}$ exists and $C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}$. The stated formula follows. $\Box$
@@
Suppose that $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ is a light weight and $D_{h}(1)_{1}=+\infty$. Then
@@
Let $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a convex light weight and $B_{h}(1)=+\infty$. Then
@@ 4 Extensions of Pólya’s theorem with $v\neq 1$
The case $v\neq 1$ when the vertex to be visited by walks differs from their starting vertex is more complicated than the previous case. The grid graph $\mathbb{Z}^{d}$ is vertex-transitive but so far this was irrelevant. For $\overline{v}\neq\overline{0}$ vertex-transitivity becomes relevant and we generalize it to $K_{\mathbb{N}}$ as follows. Let $v\in\mathbb{N}\setminus\{1\}$ with $v\neq 1$. We say that a weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is $v$-transitive if there is a bijection $f\colon\mathbb{N}\to\mathbb{N}$ such that $f(1)=v$ and $h(f[e])=h(e)$ for every $e\in\mathbb{N}_{2}$.
We again introduce some generating functions. Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}$ with $v\neq 1$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ are the walks visiting $v$ ($\neq 1$) at a step $i>0$. We are interested in the generating function
@@
The generating functions $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ are as before. We additionally introduce
@@
$E_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}e_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{e,v}(n))x^{n}\,.}$
Here $W_{b,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting and ending at $1$ and avoiding $v$. $W_{c,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding $v$. Finally, $W_{e,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding both $1$ and $v$. We have $a_{0}^{h,v}=c_{0}^{h,v}=c_{0}^{h}=e_{0}^{h,v}=0$ and $b_{0}^{h}=b_{0}^{h,v}=d_{0}^{h}=1$.
@@
Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight and the generating functions $A_{h,v}(x)$, $B_{h}(x)$, $B_{h,v}(x)$, $C_{h,v}(x)$, $D_{h}(x)$ and $E_{h,v}(x)$ in $\mathbb{C}[[x]]$ be as above. The following relations hold between them.
@@
Proof. 1. Every walk $w$ in $\bigcup_{n\geq 0}W(v,n)$ splits at the first visit of $v$ in two walks as $w=w_{1}w_{2}$. Walks $w_{1}$ are weight-counted by $C_{h,v}(x)$ and $w_{2}$ are arbitrary walks starting at $v$. By the $v$-transitivity of $h$, walks $w_{2}$ are weight-counted by $D_{h}(x)$. The first relation follows.
2. It suffices to prove the first equality. $B_{h}(x)$ weight-counts walks $w$ starting and ending at $1$. Those avoiding $v$ are weight-counted by $B_{h,v}(x)$. If $w$ visits $v$, then it uniquely splits at the first and last visits of $v$ in three walks as $w=w_{1}w_{2}w_{3}$. Walks $w_{1}$ start at $1$, end at $v$ and between avoid $v$. Walks $w_{2}$ start and end at $v$. Walks $w_{3}$ start at $v$, end at $1$ and between avoid $v$. We reverse walks $w_{3}$ and see that both $w_{1}$ and $w_{3}$ are weight-counted by $C_{h,v}(x)$. By the $v$-transitivity of $h$, the middle walks $w_{2}$ are weight-counted by $B_{h}(x)$. The second relation follows.
3. We consider walks $w$ weight-counted by $C_{h,v}(x)$. They start at $1$, end at $v$, and between avoid $v$. These walks uniquely split at the last visit of $1$ in two walks as $w=w_{1}w_{2}$. Walks $w_{1}$ are weight-counted by $B_{h,v}(x)$, and $w_{2}$ by $E_{h,v}(x)$. The third relation follows. $\Box$
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $v\in\mathbb{N}\setminus\{1\}$. If $v\not\in V(h)$ then
$a_{n}^{h,\,v}=h(W(v,\,n))=\sum_{w\in W(v,\,n)}h(w)=\sum_{w\in W(v,\,n)}0=0$}
for every $n\in\mathbb{N}_{0}$ because every walk starting at $1$ and ending at $v$ contains an edge with zero weight. Henceforth we therefore assume that $v\in V(h)$. For $w,z\in\mathbb{C}$ we define the set
@@
where if $B_{h}(1)_{1}=0$ then $B_{h,v}(1)_{1}=0$ and we interpret the fraction $\frac{0}{0}$ as $1$, so that $A_{h,v}(1)_{1}=0$.
@@
then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ and $A_{h,v}(1)_{2}$ exist, and
@@
Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed inequality hold. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.30, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ exists and $C_{h,v}(1)_{2}^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$. By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$. By item 1 of Proposition 2.25, $A_{h,v}(1)_{2}$ exists and $A_{h,v}(1)_{2}=C_{h,v}(1)_{2}D_{h}(1)_{2}$. The stated formula follows. $\Box$
@@
where if $B_{h}(1)_{1}=0$ then $B_{h,v}(1)_{1}=0$ and we interpret the fraction $\frac{0}{0}$ as $1$, so that the limit is zero.
@@
then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ and $A_{h,v}(1)_{2}$ exist, and
$\lim_{n\to\infty}a_{n}^{h,v}\in{\textstyle D_{h}(1)_{2}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
@@
Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed inequality hold. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.30, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ exists and $C_{h,v}(1)_{2}^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$. By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-x}$. Proposition 2.15 gives the stated formula:
@@
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a $v$-transitive convex light weight such that $v\in V(h)$. Suppose that $B_{h}(1)_{1}=+\infty$. Then
@@
Suppose for the contrary that $\lim_{x\to 1}F_{B_{h,v}}(x)=+\infty$. By item 3 of Proposition 4.1 and by Propositions 2.4 and 2.5 we have for every $x\in[0,1)$ the equality $F_{C_{h,v}}(x)=F_{B_{h,v}}(x)F_{E_{h,v}}(x)$. Thus $\lim_{x\to 1}F_{C_{h,v}}(x)=+\infty$. From the facts that $F_{C_{h,v}}(0)=0$, $F_{C_{h,v}}(x)$ increases and is continuous and $\lim_{x\to 1}F_{C_{h,v}}(x)=+\infty$, we obtain a (unique) number $x_{0}\in(0,1)$ such that $F_{C_{h,v}}(x)<1$ for $x\in[0,x_{0})$ and $\lim_{x\to x_{0}^{-}}F_{C_{h,v}}(x)=1$. By item 2 of Proposition 4.1 and by Propositions 2.4 and 2.5 we have for every $x\in[0,1)$ the equality

### v3

@@ 2. This follows from the inequality ($n\in\mathbb{N}_{0}$)
Proposition 2.19 for $n=2$ and item 1 of Proposition 2.25 are incomparable. The former assumes only conditional convergence, but for all three series. The latter assumes stronger absolute convergence, but only for two series. Another variant of the proposition is due to Franz (Franciszek) Mertens (1840–1927).
@@
Suppose that $U(x)$, $V(x)$ and $W(x)$ are in $\mathbb{C}[[x]]$. If $U(x)=V(x)W(x)$, and if $V(1)_{1}$ and $W(1)_{2}$ exist, then $U(1)_{1}$ exists and
@@
$S_{m}=\prod_{i=1}^{m}(v_{1}+v_{2}+\dots)=\sum_{\langle j_{1},\,\dots,\,j_{m}\rangle\in\mathbb{N}^{m}}v_{j_{1}}v_{j_{2}}\dots v_{j_{m}}\,.$}
By Proposition 2.6 it absolutely converges and has sum $(V(1)_{2}-1)^{m}$. Thus by Propositions 2.2 and 2.4, $S$ has sum $1+\sum_{m\geq 1}(-1)^{m}(V(1)_{2}-1)^{m}=\frac{1}{V(1)_{2}}$.
@@
By Proposition 2.4, $S$ has also sum $u_{0}+u_{1}+u_{2}+\cdots=U(1)_{2}$ and the stated formula follows.
@@ 2. Indeed,
By Proposition 2.4, $S$ has the sum $v_{0}+v_{1}+\dots=V(1)_{2}$. By item 1 of Proposition 2.25, $U(1)_{2}=\big(V(1)_{2}\big)^{2}$. $\Box$
@@
Finally, we consider convex light weights.
@@
We can add and remove the condition $\ell(w^{\prime})\in V(h)$ without affecting the sum because walks $w^{\prime}$ not satisfying it have zero weight. $\Box$
@@ 3 Extensions of Pólya’s theorem with $v=1$
We establish the first group of generalizations of Pólya’s theorem in Theorems 3.23.7. We consider the case $v=1$ when the vertex visited by walks coincides with their starting vertex.
We introduce some generating functions. Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ is the set of walks that visit $v$ in step $i>0$. We are interested in the generating function
$A_{h}(x)=\sum_{n\geq 0}a_{n}^{h}x^{n}=\sum_{n\geq 0}h(W(1,\,n))x^{n}$}
and define
@@
all in $\mathbb{C}[[x]]$. Here $W_{b}(n)$ are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and $W_{c}(n)$ are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and avoid $1$ between. The coefficients $a_{n}^{h}$, $\dots$, $d_{n}^{h}$ are correctly defined because they are sums of subseries of absolutely convergents series. We have $a_{0}^{h}=c_{0}^{h}=0$ and $b_{0}^{h}=d_{0}^{h}=1$.
@@ 2. We show that
$F(w)=\langle w_{1},\,w_{2},\,\dots,\,w_{j}\rangle\ \ (\in W_{c}(n_{1})\times W_{c}(n_{2})\times\dots\times W_{c}(n_{j}),\,n_{i}=m_{i}-m_{i-1})$
@@
For better orientation we label the fourteen theorems by the triples
@@
where “gen” refers to general edge weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and “con” to convex ones. The generating functions $A_{h}(x)$, $\dots$, $D_{h}(x)$ are as above.
@@
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Suppose that $D_{h}(1)_{2}$ exists and $\sum_{n\geq 1}|b_{n}^{h}|<1$. Then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $A_{h}(1)_{2}$ exists and
@@
Proof. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. By item 1 of Proposition 2.27, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h}(1)_{2}$ exists and $C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}$. By item 1 of Proposition 3.1 we have $A_{h}(x)=C_{h}(x)D_{h}(x)$. By item 1 of Proposition 2.25, $A_{h}(1)_{2}$ exists and $A_{h}(1)_{2}=C_{h}(1)_{2}D_{h}(1)_{2}$ . The stated formula follows. $\Box$
@@
Suppose that $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ is a light weight such that $D_{h}(1)_{1}=+\infty$. Then
@@
Let $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a convex light weight such that $B_{h}(1)=+\infty$. Then
@@ 4 Extensions of Pólya’s theorem with $v\neq 1$
The case when the vertex visited by walks differs from their starting vertex is more complicated. The grid graph $\mathbb{Z}^{d}$ is vertex-transitive, but so far, this has been irrelevant. For $\overline{v}\neq\overline{0}$ vertex-transitivity becomes important, and we generalize it to $K_{\mathbb{N}}$ as follows. Let $v\in\mathbb{N}\setminus\{1\}$ with $v\neq 1$. We say that a weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is $v$-transitive if there is a bijection $f\colon\mathbb{N}\to\mathbb{N}$ such that $f(1)=v$ and $h(f[e])=h(e)$ for every $e\in\mathbb{N}_{2}$.
We again introduce some generating functions. Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ are the walks visiting the vertex $v$ ($\neq 1$); necessarily in a step $i>0$. We are interested in the generating function
@@
The generating functions $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ are as before. We define
@@
$E_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}e_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{e,v}(n))x^{n}\,,}$
all in $\mathbb{C}[[x]]$. Here $W_{b,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting and ending at $1$ and avoiding $v$. $W_{c,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding $v$. Finally, $W_{e,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding both $1$ and $v$. We have $a_{0}^{h,v}=c_{0}^{h,v}=c_{0}^{h}=e_{0}^{h,v}=0$ and $b_{0}^{h}=b_{0}^{h,v}=d_{0}^{h}=1$.
@@
Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight and generating functions $A_{h,v}(x)$, $B_{h}(x)$, $B_{h,v}(x)$, $C_{h,v}(x)$, $D_{h}(x)$ and $E_{h,v}(x)$ be as above. The following relations hold between them.
@@
Proof. 1. Every walk $w$ weight-counted by $A_{h,v}(x)$ splits at the first visit of $v$ as $w=w_{1}w_{2}$. Walks $w_{1}$ are weight-counted by $C_{h,v}(x)$ and $w_{2}$ are arbitrary walks starting at $v$. By the $v$-transitivity of $h$, walks $w_{2}$ are weight-counted by $D_{h}(x)$.
2. It suffices to prove the first equality. $B_{h}(x)$ weight-counts walks $w$ starting and ending at $1$. Those avoiding $v$ are weight-counted by $B_{h,v}(x)$. If $w$ visits $v$, then it uniquely splits at the first and last visits of $v$ as $w=w_{1}w_{2}w_{3}$. Walks $w_{1}$ start at $1$, end at $v$ and between avoid $v$. Walks $w_{2}$ start and end at $v$. Walks $w_{3}$ start at $v$, end at $1$ and between avoid $v$. We reverse walks $w_{3}$ and see that both $w_{1}$ and $w_{3}$ are weight-counted by $C_{h,v}(x)$. Due to the $v$-transitivity of $h$, the middle walks $w_{2}$ are weight-counted by $B_{h}(x)$.
3. We consider walks $w$ weight-counted by $C_{h,v}(x)$. They start at $1$, end at $v$, and between avoid $v$. These walks uniquely split at the last visit of $1$ as $w=w_{1}w_{2}$. Walks $w_{1}$ are weight-counted by $B_{h,v}(x)$, and $w_{2}$ by $E_{h,v}(x)$. $\Box$
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $v\in\mathbb{N}\setminus\{1\}$. If $v\not\in V(h)$ then $a_{n}^{h,v}=\sum_{w\in W(v,n)}h(w)=\sum_{w\in W(v,n)}0=0$ for every $n\in\mathbb{N}_{0}$ because every walk starting at $1$ and visiting $v$ contains an edge with zero weight. Henceforth, we therefore assume that $v\in V(h)$. For $w,z\in\mathbb{C}$ we define the set
@@
where if $B_{h}(1)_{1}=0$ then $B_{h,v}(1)_{1}=0$ and we interpret $\frac{0}{0}$ as $1$, so that $A_{h,v}(1)_{1}=0$.
@@
then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $A_{h,v}(1)_{2}$ exists and
@@
Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.30, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ exists and $(C_{h,v}(1)_{2})^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$. By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$. By item 1 of Proposition 2.25, $A_{h,v}(1)_{2}$ exists and $A_{h,v}(1)_{2}=C_{h,v}(1)_{2}D_{h}(1)_{2}$. The stated formula follows. $\Box$
@@
where if $B_{h}(1)_{1}=0$ then $B_{h,v}(1)_{1}=0$ and we interpret $\frac{0}{0}$ as $1$, so that the limit is zero.
@@
then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$ and
$\lim_{n\to\infty}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
@@
Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.30, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ exists and $C_{h,v}(1)_{2}^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$. By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-x}$. Proposition 2.15 gives the stated formula:
@@
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a $v$-transitive convex light weight such that $v\in V(h)$. If $B_{h}(1)_{1}=+\infty$ then
@@
Suppose for the contrary that $\lim_{x\to 1}F_{B_{h,v}}(x)=+\infty$. By item 3 of Proposition 4.1 and by Propositions 2.4 and 2.5 we have for every $x\in[0,1)$ the equality $F_{C_{h,v}}(x)=F_{B_{h,v}}(x)F_{E_{h,v}}(x)$. Thus $\lim_{x\to 1}F_{C_{h,v}}(x)=+\infty$. Since $F_{C_{h,v}}(0)=0$, $F_{C_{h,v}}(x)$ increases and is continuous and $\lim_{x\to 1}F_{C_{h,v}}(x)=+\infty$, we obtain a (unique) number $x_{0}\in(0,1)$ such that $F_{C_{h,v}}(x)<1$ for $x\in[0,x_{0})$ and $\lim_{x\to x_{0}^{-}}F_{C_{h,v}}(x)=1$. By item 2 of Proposition 4.1 and by Propositions 2.4 and 2.5 we have for every $x\in[0,1)$ the equality